# From Local to Global Order: A Theory of Neural Synaptic Balance

### v1

@@ 1. Introduction
One of the most common complaints against neural networks is that they provide at best “black-box” solutions to problems. In particular, when large neural networks are trained on complex tasks, they produce large arrays of synaptic weights that have no clear structure and are difficult to interpret. Thus finding any kind of structure in the weights of large neural networks is of great interest. Here we study a particular kind of structure we call neural synaptic balance and the conditions under which it emerges. Neural synaptic balance is different from the biological notion of balance between excitation and inhibition [12, 10, 13, 17, 23]. We use the term to refer to any systematic relationship between the input and output synaptic weights of individual neurons or layers of neurons. Here we consider the case where the cost of the input weights is equal to the cost of the output weights, where the cost is defined by some regularizer. One of the most basic examples of such a relationship is when the sum of the squares of the input weights of a neuron is equal to the sum of the squares its output weights.
@@
where $IN$ and $OUT$ denote the set of incoming and outgoing weights respectively, $A=\sum_{i\in IN}w_{i}^{2}$, and $B=\sum_{i\in OUT}w_{i}^{2}$. The product of the two terms on the right-hand side of Equation 1.1 is equal to $AB$ and does not depend on $\lambda$. Thus, the minimum is achieved when these two terms are equal, which yields: $(\lambda^{*})^{4}=B/A$ for the optimal $\lambda^{*}$. The corresponding new set of weights, $v_{i}=\lambda^{*}w_{i}$ for the input weights and $v_{i}=w_{i}/\lambda^{*}$ for the outgoing weights, must be balanced: $\sum_{i\in IN}v_{i}^{2}=\sum_{i\in OUT}v_{i}^{2}$, since its optimal scaling factor can only be $\lambda=1$. Furthermore, if an entire network of ReLU neurons is properly trained using a standard error function with an $L_{2}$ regularizer, at the end of training one observes a remarkable phenomenon: for each ReLU neuron, the norm of the incoming synaptic weights is approximately equal to the norm of the outgoing synaptic weights, i.e. every neuron is balanced.
There have been isolated previous studies of this kind of synaptic balance [9, 25] under special conditions. For instance, in [9], it is shown that if a deep network is initialized in a balanced state with respect to the sum of squares metric, and if training progresses with an infinitesimal learning rate, then balance is preserved throughout training. Here, we take a different approach aimed at uncovering the generality of neuronal balance phenomena, the learning conditions under which they occur, as well as new local balancing algorithms and their convergence properties. We explain and study neural synaptic balance in its generality in terms of activation functions, regularizers, network architectures, and training stages. In particular, we systematically answer questions such as: Why does balance occur? Does it occur only with ReLU neurons? Does it occur only with $L_{2}$ regularizers? Does it occur in architectures beyond fully connected feedforward architectures (e.g. in locally connected, convolutional, or recurrent architectures)? Does it occur only at the end of training? In the process of answering these questions, we introduce local scaling and balancing operations for individual neurons or entire neural layers. Furthermore, we show that when these local operations are applied stochastically, a global balanced state always emerges, and this state is unique and depends only on the initial weights, but not on the order in which the neurons are balanced.
@@ 2.1. Additive Activation Functions
Obvisouly linear activation functions are additively linear. Conversely, if $f$ is additively linear, the following three properties are true:
@@ 2.5. BiLU Activation Functions
These include linear units ($a=b$), ReLU units ($a=0,b=1$), leaky ReLU ($a=\epsilon;b=1$) units, and symmetric linear units ($a=-b$), all of which can also be viewed as special cases of piece-wise linear units [27], with a single hinge. One advantage of ReLU and more generally BiLU neurons, which is very important during backpropagation learnin, is that their derivative is very simple and can only take one of two values ($a$ or $b$),
@@
In Appendix A, we provide a simple proof that BiLU neurons have universal approximation properties. In the next two sections, we introduce two fundamental neuronal operations, scaling and balancing, that can be applied to the incoming and outgoing synaptic weights of neurons with BiLU activation functions.
@@ 5. General Framework and Single Neuron Balance
where $W$ denotes all the weights in the network and $E(W)$ is typically the negative log-likelihood (LMS error in regression tasks, or cross-entropy error in classification tasks). We assume that the $g_{w}$ are continuous, and lower-bounded by 0. To ensure the existence and uniqueness of minimum during the balancing of any neuron, We will assume that each function $g_{w}$ depends only on the magnitude $|w|$ of the corresponding weight, and that is it continuous and monotonically increasing from 0 to +$\infty$+ ∞ ($g_{w}(0)=0$ and $\lim_{x\to+\infty}g_{w}(x)=+\infty$). Clearly, $L_{2},L_{1}$ and more generally all $L_{p}$ regularizers are special cases where, for $p>0$, $L^{p}$ regularization is defined by: $R(W)=\sum_{w}|w|^{p}$.
@@
(Balance and Regularizer Minimization) We now consider the same setting as in Theorem 5.2, but in addition we assume that the functions $g_{w}$ are continuously differentiable, except perhaps at the origin. Then, for any neuron, there exists at least one optimal value $\lambda^{*}$ that minimizes $R(W)$. This value must be a solution of the consistency equation:
@@ 6.1. Bi-Power Units (BiPU)
The set of activation functions $f$ satisfying $f(\lambda x)=\lambda^{c}f(x)$ for any $x\ in\mathbb{R}$ and any $\lambda>0$ consist of the functions of the form:
@@ 9.2. Convergence of the Basic Stochastic Balancing Algorithm to a Unique Optimum
This objective is strictly convex in the variables $L_{ij}$, as a sum of strictly convex functions (exponentials). However, to show that it is a convex optimization problem we need to study the constraints on the variables $L_{ij}$. From the set of $\Lambda_{i}$’s it is easy to construct a unique set of $L_{ij}$. However what about the converse?
@@
Note that one could coalesce all the input units and all output units into a single unit, in which case a path from an input unit to and output unit becomes also a directed cycle. In this representation, the constraints are that the sum of the $L_{ij}$ must be zero along any directed cycle. In general, it is not necessary to write a constraint for every path from input units to output units. It is sufficient to select a representative set of paths such that every unit appears in at least one path.
@@
We know that the stochastic balancing algorithm always converges to a balanced state. We need to show that it cannot converge to any other balanced state, and in fact that the global optimum is the only balanced state. By contradiction, suppose it converges to a different balanced state associated with the coordinates $(L^{B}_{ij})$ (point $B$ in Figure 6). Because of the self-consistency, this point is also associated with a unique set of $(\Lambda^{B}_{i})$ coordinates. The cost function is continuous and differentiable in both the $L_{ij}$’s and the $\Lambda_{i}$’s coordinates. If we look at the negative gradient of the regularizer, it is non-zero and therefore it must have at least one non-zero component $\partial R/\partial\Lambda_{i}$ along one of the $\Lambda_{i}$ coordinates. This implies that by scaling the corresponding unit $i$ in the network, the regularizer can be further reduced, and by balancing unit $i$ the balancing algorithm will reach a new point ($C$ in Figure 6) with lower regularizer cost. This contradicts the assumption that $B$ was associated with a balanced stated. Thus, given an initial set of weights $W$, the stochastic balancing algorithm must always converge to the same and unique optimal balanced state $W^{*}$ associated with the self-consistent point $A$. A particular stochastic schedule corresponds to a random path within the linear manifold from the origin (at time zero all the multipliers are equal to 1, and therefore $M_{ij}=1$ and $L_{ij}=0$) for any $i$ and any $j$ to the unique optimum point $A$. ∎
@@
It should be clear from the proof that the same result holds also for any deterministic balancing schedule, as well as for tied and non-tied subset balancing, e.g., for layer-wise balancing and tied layer-wise balancing. In the Appendix, we provide an analytical solution for the case of tied layer-wise balancing in a layered feed-forward network.
@@
It should be clear from the proof that the same convergence to the unique global optimum is observed if each neuron, when stochastically visited, is favorably scaled rather than balanced, i.e., it is scaled with a factor that reduces $R$ but not necessarily minimizes $R$. Stochastic balancing can also be viewed as a form of EM algorithm where the E and M steps can be taken fully or partially.
Figure 7. SGD applied to $E$ alone, in general, does not converge to a balanced state, but sGD applied to $E+R$ converges to a balanced state. (A-C) Simulations use a deep fully connected autoencoder trained on the MNIST dataset. (D-F) Simulations use deep locally connected network trained on the CFAR10 dataset. (A,D) Regularization leads to neural balance. (B,E) The training loss decreases and converges during training (these panels are not meant for assessing the quality of learning when using a regularizer). (C,F) Using weight regularization decreases the norm of weights. (A-F) Shaded areas correspond to one s.t.d around the mean (in some cases the s.t.d. is small and the shaded area is not visible).
@@ 12. Conclusion
The theory of neural synaptic balance explains some basic findings regarding $L_{2}$ balance in feedforward networks of ReLU neurons and extends them in several directions. The first direction is the extension to BiLU and other activation functions (BiPU). The second direction is the extension to more general regularizers, including all $L_{p}$ ($p>0$) regularizers. The third direction is the extension to non-layered architectures, recurrent architectures, convolutional architectures, as well as architectures with mixed activation functions. The theory is based on two local neuronal operations: scaling which is commutative, and balancing which is not commutative. Finally, and most importantly, given any initial set of weights, when local balancing operations are applied in a stochastic or deterministic manner, global order always emerges through the convergence of the balancing algorithm to the same unique set of balanced weights. The reason for this convergence is the existence of an underlying convex optimization problem where the relevant variables are constrained to a linear, only architecture-dependent, manifold. Scaling and balancing operations are local and thus may have applications in physical, non-digitally simulated, neural networks.

### v2

@@ 1. Introduction
One of the most common complaints against neural networks is that they provide at best “black-box” solutions to problems. In particular, when large neural networks are trained on complex tasks, they produce large arrays of synaptic weights that have no clear structure and are difficult to interpret. Thus finding any kind of structure in the weights of large neural networks is of great interest. Here we study a particular kind of structure we call neural synaptic balance and the conditions under which it emerges. Neural synaptic balance is different from the biological notion of balance between excitation and inhibition [12, 10, 13, 17, 23]. We use the term to refer to any systematic relationship between the input and output synaptic weights of individual neurons or layers of neurons. Here we consider the case where the cost of the input weights is equal to the cost of the output weights, where the cost is defined by some regularizer. One of the most basic examples of such a relationship is when the sum of the squares of the input weights of a neuron is equal to the sum of the squares of its output weights.
@@
where $IN$ and $OUT$ denote the set of incoming and outgoing weights respectively, $A=\sum_{i\in IN}w_{i}^{2}$, and $B=\sum_{i\in OUT}w_{i}^{2}$. The product of the two terms on the right-hand side of Equation 1.1 is equal to $AB$ and does not depend on $\lambda$. Thus, the minimum is achieved when these two terms are equal, which yields: $(\lambda^{*})^{4}=B/A$ for the optimal $\lambda^{*}$. The corresponding new set of weights, $v_{i}=\lambda^{*}w_{i}$ for the input weights and $v_{i}=w_{i}/\lambda^{*}$ for the outgoing weights, must be balanced: $\sum_{i\in IN}v_{i}^{2}=\sum_{i\in OUT}v_{i}^{2}$. This is because the optimal scaling factor for these rescaled weights must be $\lambda^{*}=1$. Furthermore, if an entire network of ReLU neurons is properly trained using a standard error function with an $L_{2}$ regularizer, at the end of training one observes a remarkable phenomenon: for each ReLU neuron, the norm of the incoming synaptic weights is approximately equal to the norm of the outgoing synaptic weights, i.e. every neuron is balanced.
There have been isolated previous studies of this kind of synaptic balance [9, 25] under special conditions. For instance, in [9], it is shown that if a deep network is initialized in a balanced state with respect to the sum of squares metric, and if training progresses with an infinitesimal learning rate, then balance is preserved throughout training. Here, we take a different approach aimed at uncovering the generality of neuronal balance phenomena, the learning conditions under which they occur, as well as new local balancing algorithms and their convergence properties. We explain and study neural synaptic balance in its generality in terms of activation functions, regularizers, network architectures, and training stages. In particular, we systematically answer questions such as: Why does balance occur? Does it occur only with ReLU neurons? Does it occur only with $L_{2}$ regularizers? Does it occur only in fully connected feedforward architectures (as opposed to, for instance, locally connected, convolutional, or recurrent architectures)? Does it occur only at the end of training? In the process of answering these questions, we introduce local scaling and balancing operations for individual neurons or entire neural layers. Furthermore, we show that when these local operations are applied stochastically, a global balanced state always emerges, and this state is unique and depends only on the initial weights, but not on the order in which the neurons are balanced.
@@ 2.1. Additive Activation Functions
Obviously linear activation functions are additively linear. Conversely, if $f$ is additively linear, the following three properties are true:
@@ 2.5. BiLU Activation Functions
These include linear units ($a=b$), ReLU units ($a=0,b=1$), leaky ReLU ($a=\epsilon;b=1$) units, and symmetric linear units ($a=-b$), all of which can also be viewed as special cases of piece-wise linear units [27], with a single hinge. One advantage of ReLU and more generally BiLU neurons, which is very important during backpropagation learning, is that their derivative is very simple and can only take one of two values ($a$ or $b$).
@@
In Appendix A, we provide a simple proof that networks of BiLU neurons, even with a single hidden layer, have universal approximation properties. In the next two sections, we introduce two fundamental neuronal operations, scaling and balancing, that can be applied to the incoming and outgoing synaptic weights of neurons with BiLU activation functions.
@@ 5. General Framework and Single Neuron Balance
where $W$ denotes all the weights in the network and $E(W)$ is typically the negative log-likelihood (LMS error in regression tasks, or cross-entropy error in classification tasks). We assume that the $g_{w}$ are continuous, and lower-bounded by 0. To ensure the existence and uniqueness of minimum during the balancing of any neuron, We will assume that each function $g_{w}$ depends only on the magnitude $|w|$ of the corresponding weight, and that $g_{w}$ is monotonically increasing from 0 to +$\infty$+ ∞ ($g_{w}(0)=0$ and $\lim_{x\to+\infty}g_{w}(x)=+\infty$). Clearly, $L_{2},L_{1}$ and more generally all $L_{p}$ regularizers are special cases where, for $p>0$, $L^{p}$ regularization is defined by: $R(W)=\sum_{w}|w|^{p}$.
@@
(Balance and Regularizer Minimization) We now consider the same setting as in Theorem 5.2, but in addition we assume that the functions $g_{w}$ are continuously differentiable, except perhaps at the origin. Then, for any neuron, there exists at least one optimal value $\lambda^{*}$ that minimizes $R(W)$. Any optimal value must be a solution of the consistency equation:
@@ 6.1. Bi-Power Units (BiPU)
The set of activation functions $f$ satisfying $f(\lambda x)=\lambda^{c}f(x)$ for any $x\ \in\mathbb{R}$ and any $\lambda>0$ consist of the functions of the form:
@@ 9.2. Convergence of the Basic Stochastic Balancing Algorithm to a Unique Optimum
This objective is strictly convex in the variables $L_{ij}$, as a sum of strictly convex functions (exponentials). However, to show that it is a convex optimization problem we need to study the constraints on the variables $L_{ij}$. In particular, from the set of $\Lambda_{i}$’s it is easy to construct a unique set of $L_{ij}$. However what about the converse?
@@
One could coalesce all the input units and all output units into a single unit, in which case a path from an input unit to and output unit becomes also a directed cycle. In this representation, the constraints are that the sum of the $L_{ij}$ must be zero along any directed cycle. In general, it is not necessary to write a constraint for every path from input units to output units. It is sufficient to select a representative set of paths such that every unit appears in at least one path.
@@
We know that the stochastic balancing algorithm always converges to a balanced state. We need to show that it cannot converge to any other balanced state, and in fact that the global optimum is the only balanced state. By contradiction, suppose it converges to a different balanced state associated with the coordinates $(L^{B}_{ij})$ (point $B$ in Figure 6). Because of the self-consistency, this point is also associated with a unique set of $(\Lambda^{B}_{i})$ coordinates. The cost function is continuous and differentiable in both the $L_{ij}$’s and the $\Lambda_{i}$’s coordinates. If we look at the negative gradient of the regularizer, it is non-zero and therefore it must have at least one non-zero component $\partial R/\partial\Lambda_{i}$ along one of the $\Lambda_{i}$ coordinates. This implies that by scaling the corresponding unit $i$ in the network, the regularizer can be further reduced, and by balancing unit $i$ the balancing algorithm will reach a new point ($C$ in Figure 6) with lower regularizer cost. This contradicts the assumption that $B$ was associated with a balanced stated. Thus, given an initial set of weights $W$, the stochastic balancing algorithm must always converge to the same and unique optimal balanced state $W^{*}$ associated with the self-consistent point $A$. A particular stochastic schedule corresponds to a random path within the linear manifold from the origin (at time zero all the multipliers are equal to 1, and therefore for any $i$ and any $j$: $M_{ij}=1$ and $L_{ij}=0$) to the unique optimum point $A$. ∎
@@
From the proof, it is clear that the same result holds also for any deterministic balancing schedule, as well as for tied and non-tied subset balancing, e.g., for layer-wise balancing and tied layer-wise balancing. In the Appendix, we provide an analytical solution for the case of tied layer-wise balancing in a layered feed-forward network.
@@
The same convergence to the unique global optimum is observed if each neuron, when stochastically visited, is favorably scaled rather than balanced, i.e., it is scaled with a factor that reduces $R$ but not necessarily minimizes $R$. Stochastic balancing can also be viewed as a form of EM algorithm where the E and M steps can be taken fully or partially.
Figure 7. SGD applied to $E$ alone, in general, does not converge to a balanced state, but sGD applied to $E+R$ converges to a balanced state. (A-C) Simulations use a deep fully connected autoencoder trained on the MNIST dataset. (D-F) Simulations use a deep locally connected network trained on the CFAR10 dataset. (A,D) Regularization leads to neural balance. (B,E) The training loss decreases and converges during training (these panels are not meant for assessing the quality of learning when using a regularizer). (C,F) Using weight regularization decreases the norm of weights. (A-F) Shaded areas correspond to one s.t.d around the mean (in some cases the s.t.d. is small and the shaded area is not visible).
@@ 12. Conclusion
The theory of neural synaptic balance explains some basic findings regarding $L_{2}$ balance in feedforward networks of ReLU neurons and extends them in several directions. The first direction is the extension to BiLU and other activation functions (BiPU). The second direction is the extension to more general regularizers, including all $L_{p}$ ($p>0$) regularizers. The third direction is the extension to non-layered architectures, recurrent architectures, convolutional architectures, as well as architectures with mixed activation functions. The theory is based on two local neuronal operations: scaling which is commutative, and balancing which is not commutative. Finally, and most importantly, given any initial set of weights, when local balancing operations are applied in a stochastic or deterministic manner, global order always emerges through the convergence of the balancing algorithm to the same unique set of balanced weights. The reason for this convergence is the existence of an underlying convex optimization problem where the relevant variables are constrained to a linear, only architecture-dependent, manifold. Scaling and balancing operations are local and thus may have applications in physical, non-digitally simulated, neural networks where the emergence of global order from local operations may lead to better operating characteristics and lower energy consumption.